# Investigating Feature and Model Importance in Android Malware Detection: An Implemented Survey and Experimental Comparison of ML-Based Methods

### v3

@@
Abstract.
The popularity of Android means it is a common target for malware. Over the years, various studies have found that machine learning models can effectively discriminate malware from benign applications. However, as the operating system evolves, so does malware, bringing into question the findings of these previous studies, many of which report very high accuracies using small, outdated, and often imbalanced datasets. In this paper, we reimplement 18 representative past works and reevaluate them using a balanced, relevant, and up-to-date dataset comprising 124,000 applications. We also carry out new experiments designed to fill holes in existing knowledge, and use our findings to identify the most effective features and models to use for Android malware detection within a contemporary environment. We show that high detection accuracies (up to 96.8%) can be achieved using features extracted through static analysis alone, yielding a modest benefit (1%) from using far more expensive dynamic analysis. API calls and opcodes are the most productive static and TCP network traffic provide the most predictive dynamic features. Random forests are generally the most effective model, outperforming more complex deep learning approaches. Whilst directly combining static and dynamic features is generally ineffective, ensembling models separately leads to performances comparable to the best models but using less brittle features.
1. Introduction
@@
Smartphone usage has increased exponentially over recent years. According to Ericsson’s mobility report (Ericsson, 2023), there were 6.9 billion active smartphone subscriptions in 2023, with 70.48% of them using an Android smartphone (StatCounter, 2024). It is common practice to augment smartphones with applications, which in Android can be downloaded from the official Google Play Store (Play, 2024) or third-party application stores. The availability and easy access to applications through third party stores, in particular, provides attackers with a means of distributing malware. According to the latest report published by the computer security company G Data, a new piece of Android malware appears on the Internet every 12 seconds (DATA, 2022).
@@
Several approaches have been proposed to secure mobile operating systems (OS), including the application sandbox approach used by Android (Elenkov, 2015). In order to grant access to device services outside of the sandbox, Android uses a permissions system; however, this has its own shortcomings (Almomani and Khayer, 2020). A number of security solutions, including malware detectors, vulnerability detection, user and developer reviews have been proposed (Sufatrio et al., 2015). Of these solutions, Android malware detection, or anti-malware, is one of the most widely used. Anti-malware can initially be used to stop an application from being released into application stores or later at the user level, preventing the user from installing the application.
Traditional non-machine learning approaches to malware detection use signatures to detect malicious files. However, any slight variation to the file might cause the anti-malware to not detect the malware. In particular, this makes it difficult to detect zero-day attacks or existing malware variations. With the rapid growth of Android malware, there may be occasions when signature-based anti-malware will not be able to detect thousands, and potentially more, existing malware. Machine learning (ML) can be used to solve this problem. ML-based techniques require analysis of Android applications to extract features that represent the characteristics of the application. The analysis might be static or dynamic. Static analysis extracts features from the source code of the application, whereas dynamic analysis extracts features observed whilst monitoring the behaviour of the application in a running state (Sikorski and Honig, 2012).
We recently surveyed existing approaches that use machine learning to detect Android malware (Muzaffar et al., 2022). One of the insights from our study was that most of these approaches, including many of the recent ones, were evaluated using obsolete versions of Android running historic, and often quite small, collections of Android applications. This observation has also been made in a recent commentary piece (Guerra-Manzanares, 2023). This makes it difficult to determine both whether these approaches will work in the current Android ecosystem, and whether they will generalise to larger collections of Android applications. Since they were evaluated using a variety of different metrics on many different datasets, it is also difficult to infer meaningful comparisons regarding the importance of feature and model choices.
@@
To address these problems, the aim of this study is to rigorously analyse the importance of feature and model choices while training ML models for Android malware detection. We primarily do this by reimplementing past works and reevaluating them using a large dataset that comprises current Android applications that we collected during the period of 2019–2021, allowing us to carry out a meaningful comparison of these methods within the current Android ecosystem and on a level playing-field. It is not feasible, nor perhaps desirable, to reimplement all the previous work in this area. Instead we focus on a group of studies which are representative of the diversity of modelling approaches used in this field, favouring those which have reported high levels of malware discrimination. This choice was guided by our previous review of published work on ML-based Android malware detection (Muzaffar et al., 2022). We also conduct a number of new experiments to fill in knowledge gaps in the existing literature, and consider the benefits of using ensemble models that combine existing approaches.
@@
We review the Android feature space by reimplementing past works. To the best of our knowledge, this is the first-of-its-kind review and analysis of the importance of different static and dynamic features while training ML models.
@@
We compile a balanced up-to-date dataset of Android malware and benign applications that is, at the time of writing this paper, the largest publicly available dataset for assessing Android anti-malware. We also share the tools used to create this dataset to make it easier for the community to develop and assess future anti-malware.
Using this dataset, we reimplement and reevaluate past Android anti-malware approaches that used static, dynamic and hybrid feature analysis approaches.
We rigorously compare different types of ML models to determine which ones work best with particular Android feature sets.
We report the best performing features, models, and feature selection algorithms for Android malware detection.
@@
We present an ensemble approach that leverages the best performing static and dynamic models, achieving an accuracy of 97.8% on our contemporary dataset.
@@
The paper is organized as follows. Section 2 describes our methodology, including dataset collection, our static and dynamic analysis tools, and our evaluation framework. This methodology is then applied to static analysis, dynamic analysis, and hybrid analysis — the three main branches of feature analysis and model building in malware detection — in Sections 3, 4 and 5, respectively. Section 7 then considers ensemble methods that combine the best models from the previous sections. Section 8 discusses the main findings of the study, Section 9 discusses its limitations and future work, and Section 10 concludes.
2. Methodology
@@
In this section, we present our framework for assessing Android anti-malware approaches. Section 2.1 describes the dataset that we use and how it was collected. Section 2.2 describes the tool we developed for performing static and dynamic analysis of Android applications. Section 2.3 outlines the metrics we use to assess ML models. Section 2.4 motivates the choice of features and our focus on binary classification. Section 2.5 presents a core set of ML models and feature selection methods which we use throughout the study.
2.1. Dataset Collection
@@
We collected111Dataset available at https://researchportal.hw.ac.uk/en/datasets/android-dataset-for-malware-detection Android applications released from 2019-2021 from the stores and repositories which are typically used by end users. The lack of availability of scripts to build datasets prompted us to build crawlers to download these applications. Through periodic use, these platform-independent Python scripts can be used to maintain an up-to-date dataset. The stores we targeted were UpToDown (Uptodown, 2024), APKMirror (APKMirror, 2024) and F-Droid (F-Droid, 2024). The scripts crawl the websites for all applications irrespective of the categories the applications are filed under. We downloaded a total of 62,000 benign applications. To the best of our knowledge, this is the most realistic and up to date benign dataset available right now.
@@
To collect malware, we used the most recent Android malware dataset from VirusShare (VirusShare.com, 2024) for malware released in 2020. We also identified malware while downloading applications from the application stores. We used VirusTotal (VirusTotal, 2024) reports to label all the applications. To prevent false positives in the benign dataset, we only included applications with zero positive tags from anti-malware in VirusTotal reports. The final malware dataset consisted of 62,000 applications, matching the size of the benign dataset.
2.2. DroidDissector
We developed an automated analysis tool, DroidDissector222Code and scripts available at https://github.com/alisolehria/DroidDissector-A-Static-and-Dynamic-Analysis-Tool-for-Android-Malware-Detection, to extract both static and dynamic features from Android applications. Drawing from the most prominent features highlighted in the literature, our tool attempts to provide coverage of the whole Android malware detection landscape (Muzaffar et al., 2023). The tool’s architecture is depicted in Figure 1. It comprises two sub-systems: one for extracting static features, and another for extracting dynamic features.
2.2.1. Static Analysis Sub-System
APK files are first reverse-engineered and decompressed using APKtool (Connor Tumbleson, 2019) to decode the manifest file and smali files from the APK package. Permissions used by the application, the application package name and the different application components — activities, services, broadcast receivers, and content providers — are stated in the Android manifest file. APKtool decodes the dex (binary code) files into readable Java-like smali files. APKTool creates a directory for each APK file. The static analysis tool then analyzes the files produced by the APKTool to retrieve features. For each application, the features extracted by our static analysis tool include:
- Hardware Components::
- Requested Permissions::
- App Components::
- Filtered Intents::
- Used Permissions::
- Network Addresses::
- API Calls v30::
- Restricted API Calls::
List of restricted API calls defined by Drebin (Arp et al., 2014). We built this feature set by analyzing the smali files extracted from the APK.
- Suspicious API Calls::
List of suspicious API calls defined by Drebin (Arp et al., 2014). We built this feature set by analyzing the smali files extracted from the APK.
- API Call Graphs::
- Opcodes::
2.2.2. Dynamic Analysis Sub-System
The dynamic analysis sub-system runs on the readily-available Android emulators provided by Android SDK (Android Developers, 2024a). The virtual images in the emulators also allow root access, which is essential for dynamic analysis. Here we describe the environment required to run the tool and the features extracted by the tool.
2.2.3. Environment
Python 3: All the dynamic modules require Python to run.
Android Emulator: Our tool uses the emulator provided by Android SDK as the virtual environment to run the analysis. The emulator should be on the system path. We have tested the tool on Android 8, 9, 10 and 11 and we expect it to run on the upcoming Android 12.
2.2.4. Feature Extraction
- System Calls::
- Network Traffic::
Our tool uses Android emulator’s tcpdump to collect the complete network traffic of the application. It stores the resulting data in “pcap” files, later used to analyze and extract features. Wireshark (Wirshark, 2024) can process this file to analyze the application’s network traffic. It also saves a separate network traffic file for every application.
- Android Logcat::
- API Calls::
The tool requires hooking of API calls beforehand to monitor execution of API calls during the run-time of an application. However, it is not computationally feasible to monitor the 130,000+ API calls present in the Android SDK. Therefore, the tool carries out the following steps: before starting the dynamic analysis, it checks if a corresponding static analysis report is present for the application. In addition to hooking the classic calls used in the literature of dynamic analysis tools (MobSF, 2023), we also hooked application-specific API calls. If present, it hooks the API calls extracted in the static analysis report, or else, it performs a static analysis on the application using the static analysis tool. The dynamic analysis tool then hooks the API calls to monitor during the execution of the application using Frida. The tool saves a separate file with all the APIs called during execution for further analysis. As the process may involve static analysis, the API call module is optional in the dynamic analysis tool. By default, the tool monitors a pre-defined list of API calls; however to monitor the API calls mined from an application’s static analysis, the static analysis module is required.
2.3. Evaluation Metrics

### v4

# Reassessing feature-based Android malware detection in a contemporary context

@@
Reassessing feature-based Android malware detection in a contemporary context
Ali Muzaffar*1, Hani Ragab Hassen1, Hind Zantout1, Michael A Lones2,
1 Heriot-Watt University, Dubai, UAE
2 Heriot-Watt University, Edinburgh EH14 4AS, United Kingdom
* ali.muzaffar@hw.ac.uk
1 Abstract
We report the findings of a reimplementation of 18 foundational studies in feature-based machine learning for Android malware detection, published during the period 2013–2023. These studies are reevaluated on a level playing field using a contemporary Android environment and a balanced dataset of 124,000 applications. Our findings show that feature-based approaches can still achieve detection accuracies beyond 98%, despite a considerable increase in the size of the underlying Android feature sets. We observe that features derived through dynamic analysis yield only a small benefit over those derived from static analysis, and that simpler models often out-perform more complex models. We also find that API calls and opcodes are the most productive static features within our evaluation context, network traffic is the most predictive dynamic feature, and that ensemble models provide an efficient means of combining models trained on static and dynamic features. Together, these findings suggest that simple, fast machine learning approaches can still be an effective basis for malware detection, despite the increasing focus on slower, more expensive machine learning models in the literature.
2 Introduction
@@
Smartphone usage has increased exponentially over recent years. According to Ericsson’s mobility report [ericsson2023], there was a year-on-year increase of 6 per cent in the number of smartphones in the first quarter of 2024, with 71.67% of them using an Android smartphone [androidData]. It is common practice to augment smartphones with applications, which in Android can be downloaded from the official Google Play Store [playstore] or third-party application stores. The availability and easy access to applications through third party stores, in particular, provides attackers with a means of distributing malware. According to the latest report published by the computer security company G Data, a new piece of Android malware appears on the Internet every 12 seconds [malwareStat].
@@
Various methods have been proposed to secure mobile operating systems (OS), including the application sandbox approach used by Android [Elenkov2015]. In order to grant access to device services outside of the sandbox, Android uses a permissions system; however, this has its own shortcomings [Almomani2020]. A number of security solutions, including malware detectors, vulnerability detection, user and developer reviews have been proposed [Sufatrio2015]. Of these solutions, Android malware detection, or anti-malware, is one of the most widely used. Anti-malware can initially be used to stop an application from being released into application stores or later at the user level, preventing the user from installing the application. Traditional approaches to malware detection use signatures to detect malicious files. However, any slight variation to the file might cause this approach to fail. In particular, this makes it difficult to detect zero-day attacks or existing malware variations. With the rapid growth of Android malware, there may be occasions when signature-based anti-malware will not be able to detect thousands, and potentially more, existing malware. Machine learning (ML) based approaches, on the other hand, attempt to get around these problems by learning models of malware that generalise beyond single applications. These models can then be used to detect malware samples that were not seen during training.
ML is a fast-changing field. One recent change is the move from feature-based approaches to end-to-end learning. In the context of Android malware detection, feature-based approaches involve analysing applications in order to extract features that can then be used to train models. This analysis might be static or dynamic. Static analysis uses domain knowledge to extract relevant features from the source code of the application, whereas dynamic analysis extracts features observed whilst monitoring the behaviour of the application in a running state [Sikorski2012]. End-to-end approaches, by comparison, do not delineate feature extraction from modelling. Instead, features are implicitly learnt as part of the model training process. In theory, this can lead to more sensitive modelling of patterns in data, but it comes with greater computational expense, lower interpretability, and an increased likelihood of overfitting.
In 2022, we surveyed approaches that use ML to detect Android malware [Muzaffar2022], and observed that most feature-based approaches report detection accuracies well above 95%, in some cases approaching 100%. Since then, the focus of new studies has moved on to more complex end-to-end modelling approaches such as transformers. However, given that published accuracies were already very high, this begs the question of whether there is a need for more complex, more expensive, approaches.
As discussed by us in [Muzaffar2022], and more recently by Guerra-Manzanares [guerra2023android], published accuracy figures can be unreliable. There are several reasons for this. One is the historical context in which they were evaluated; Android malware continues to evolve, and approaches that worked in the past may not work today. Further to this are weaknesses in the original evaluation methodologies; many were evaluated with obsolete versions of Android running historic, and often quite small datasets. Third, ML as a modelling approach has considerable problems in terms of replicability [lones2024avoiding, kapoor2024reforms]. It is common for the results of reimplemented studies to be significantly different to the originally published results [KAPOOR2023100804, gibney2022ai]. This is often due to errors made by practitioners, and it has been reported that such errors are commonplace in the security domain [arp2022and].
@@
Within this context, the aim of this study is to determine whether feature-based approaches to Android malware detection are still sufficient to detect malware within a contemporary Android environment. We primarily do this by reimplementing foundational works from the period 2013-2023 and reevaluating them using a large dataset that comprises recent Android applications collected during the period 2019–2021. It is not feasible, nor perhaps desirable, to reimplement all the previous work in this area. Instead we focus on a group of studies which are representative of the diversity of modelling approaches used in this field, favouring those which have reported high levels of malware discrimination, guided by [Muzaffar2022]. Our emphasis is on the comparative analysis of feature representations, model selection and feature selection strategies rather than on the chronological progression of research. We also conduct a number of new experiments to fill in knowledge gaps in the existing literature, and consider the benefits of using ensemble models that combine existing approaches.
@@
We reimplement and reevaluate 18 foundational studies in feature-based Android anti-malware, and find that the majority still achieve accuracies beyond 95%, with the best models reaching above 98%. This questions the increasing focus on more complex and more expensive models in malware detection.
@@
We assess these approaches on a level playing field using a balanced and up-to-date dataset of Android malware and benign applications. We share both this dataset and the tools used to create it to make it easier for the community to develop and evaluate future anti-malware.
We find that ML models with static features often perform well against those with more expensive dynamic and hybrid features. We also find that more traditional ML models often out-perform deep learning models, which again questions the increasing focus on more complex approaches.
@@
We present an ensemble model that leverages the best performing static and dynamic models, achieving the highest accuracy and true positive rate on our contemporary dataset. Notably this does not require the use of brittle network features, which are used by the best performing individual models.
We study the role of feature selection, and show that significantly better models can be produced through aggressive feature selection. This also provides a route to handle challenges caused by the rapidly increasing complexity of Android, which has considerably boosted feature counts.
@@
The paper is organized as follows. Section 4 describes our methodology, including dataset collection, feature extraction, and our evaluation framework. This methodology is then applied to static analysis, dynamic analysis, and hybrid analysis — the three main branches of feature analysis and model building in malware detection — in Sections 5, 6 and 7, respectively. Section 8 then presents the best individual models from the previous sections and uses ensemble methods to combine these. Section 9 discusses the main findings of the study, Section 10 discusses limitations and future work, and Section 11 concludes.
3 Background
There is an extensive literature on the use of ML in Android malware detection, and many of these past studies have reported a strong ability for ML models to discriminate malware from benign applications. In general, ML models are not trained directly on raw application files. Rather, static analysis and dynamic analysis are used to extract features that are then used to train and evaluate models. To place our study in context, this section provides a brief review of past works on Android malware detection based on ML, focusing on the broad directions of travel. For a more in depth review of the literature in this area, we point readers to [Muzaffar2022].
3.1 Static approaches
Android applications are distributed in APK (Android Package Kit) format, which can be unpacked into metadata and source code files. Static analysis is the process of extracting relevant information from these files. It does not involve any information about the application’s runtime behaviour.
Permissions are a core part of the Android security model and have shown strong performance when used as the primary feature in ML models. Permissions represent a relatively concise feature set, with up to 166 permissions depending on the Android OS version. Peiravian and Zhu [Peiravian2013] were one of the earliest to use them to train ML models, and reported promising results. Using various feature selection approaches, later studies then narrowed the range of permissions used. These reported accuracies of up to 99.6% [Wang2014, Rathore2021, Sahin2023, Ojo2023, mahindru2024permdroid], though it is surprising that such high accuracies can be achieved using features which do not directly relate to an application’s source code.
Another important source of features in previous studies has been the use of API calls. Unlike the use of permissions, this directly captures properties of an application’s source code. The simplest approach captures whether each API call is used, or how often it is used. More complex approaches involve analysing the sequence of API calls or the structure of the API call graph. ML models trained on API call features seem to perform as well as, or in some cases better than, those that use permissions [Ma2019, muzaffar2023android, wang2023android], especially when feature selection is used [Jung2018, Muzaffar2021]. However, API call information can be challenging to work with in practice because of the large number of different API calls that may occur in source code. For instance, there are currently over 100,000 distinct calls available in the Android API.
Another reportedly productive approach has been the use of features that represent Android opcodes, the low-level virtual machine instructions into which source code is compiled. Past works have represented opcodes as n-grams [Kang2016], or used natural language processing representations such as word2vec [yeboah2022], reporting F1 scores of up to 98%. Opcode sequences have also been translated into psuedoimages for use with convolutional neural networks (CNNs) [Xiao2019a, daoudi2021dexray].
Another influential feature set is the one first used within Arp et al’s highly-cited 2014 tool Drebin [Arp2014], where accuracies of up to 94% were reported. This comprised a more diverse range of static features, including certain permissions and API calls in addition to other application components. Despite its age, more recent studies have continued to use either the Drebin feature set or its accompanying application data set [Muzaffar2022].
The last few years have seen the use of increasingly more complex deep learning models to identify malware, reflecting broader trends towards increasing complexity within ML [zhao2025android, sun2025malicious]. However, it is notable that the performance metrics reported in these studies have not increased significantly above those reported in earlier studies, and this places some questions on the need for more complex approaches.
3.2 Dynamic and hybrid approaches
Dynamic analysis involves the extraction of information at runtime, i.e. whilst an Android application is executing. This involves running the application on a virtual or real device and monitoring its behaviour. Dynamic analysis has the potential to provide more relevant features for malware detection. For example, API call features can also be extracted during dynamic analysis, and when done at this stage, give an indication of actual call usage rather than potential API calls [Afonso2015]. However, this comes at the expense of a more complex and resource-intensive extraction process.
The added complexity of dynamic analysis means that there are fewer studies that involve dynamic features. However, one feature which has been widely used in these studies is system calls, which are low-level requests made to the operating system kernel. System calls, like API calls, can be represented as sequences [Malik2019] or using numeric features that capture frequency or usage [Ananya2020]. Accuracies of up to 99.4% have been reported for ML models that use system call features.
Network traffic generated while an application is running is another dynamic feature which appears to be effective for identifying malware, with accuracies of over 98% consistently reported in past works [Zulkifli2018, Wang2016]. However, in practice, there are challenges around collecting reliable network traces, especially when dynamic analysis is done in a virtual environment.
Some past works have used a combination of static and dynamic features to build ML models. Kandukuru and Sharma [Kandukuru2017] and Shyong et al. [Shyong2020] both combined permissions with network traffic features, again reporting high accuracies. Other combinations, including permissions and system calls [Kapratwar2017], have also shown promising results.
3.3 End-to-end approaches
End-to-end approaches learn the entire process without intervention, from input data to classifying malware or benign samples. End-to-end approaches typically use transformers, process low-level data (e.g. sequences of API calls or raw code), and do not explicitly carry out feature extraction or feature selection. The benefit of this is the need for less manual feature engineering and more streamlined pipelines that can lead to easier deployments. However, the downside is that the models may learn meaningless features, and they are typically expensive and slow to use in comparison to feature-based ML.
The published results of studies suggest that end-to-end approaches do not currently outperform feature-based ML-based approaches. Chimezie Obidiagha et al. [Obidiagha2024] use payloads directly as input and feed them into transformers, achieving accuracy rates of up to 96%. Souani et al.[Souani2022] utilised a BERT model on Android permissions and achieved 97% accuracy in the binary classification of apps. Bourebaa and Benmohammed [Bourebaa2025] utilised permissions and API calls with a range of more efficient BERT transformer models, reporting accuracies of up to 91.6%. Alshomrani et al. [alshomrani2024survey] survey other recent approaches.
These results suggest that, despite the ease and appeal of end-to-end learning, such approaches may not offer significant advantages in terms of accuracy over traditional ML methods, making them potentially less beneficial for practical deployments. However, it should be borne in mind that the performance metrics published in feature-based studies may not accurately reflect their contemporary utility, and this further justifies our aim of reimplementing these studies, in order to give a more realistic baseline against which to compare more recent approaches.
3.4 Challenges
The performance metrics reported in the literature might suggest that Android malware detection is largely a solved problem, yet there are a number of reasons to doubt whether these metrics are reliable indicators of how well these approaches would work in practice [Muzaffar2022, guerra2024machine].
Most notably, many of the studies cited above used small, dated or inbalanced collections of Android malware and benign applications [Muzaffar2022] This includes the 2014 Drebin dataset [Arp2014] and the 2017 AMD dataset [AMD]. In the years (and decades) since these datasets were released, malware has changed significantly, so the conclusions drawn by these studies may be unreliable.
In addition, most studies used historic versions of the Android operating system; often these were considered historic at the time of the study, with a lack of up-to-date feature extraction software preventing the use of a more contemporary environment [Muzaffar2022]. This means that the feature sets used are not representative of current Android versions; for example, the number of API calls has grown from a little over 1000 to well over 100,000 since the earliest studies cited above were published.
There are also issues with reporting standards within the Android malware detection literature. Imbalanced datasets are commonly used, yet some studies only report accuracy values, which can be very misleading when derived from imbalanced data [lones2024avoiding]. Other studies report a limited range of metrics which do not enable a full understanding of their ML model’s behaviour. Despite the widespread use of stochastic models, many studies only report the performance of a single model, or provide averages without measures of spread. Furthermore, statistical tests are rarely used to compare models, which further limits the reliability of findings.
These issues, combined with the general high error rates in ML-based studies [KAPOOR2023100804], leaves the community in a position where it is challenging to draw robust conclusions from the published literature. For instance, many studies have found that relatively simple features in combination with relatively simple ML models can be used to consistently identify malware. It is unclear whether this remains the case today and it is important to establish the current state of play given the recent proliferation of AI genrated malware and to understand whether there is any need for more complex approaches.
The primary purpose of this study is to cast new light on the existing literature, by reavaluating existing approaches within a contemporary Android environment, using a large balanced up-to-date dataset, and following rigourous reporting standards. We also extend the existing literature in a number of ways by filling in notable gaps and by applying ensemble methods that combine existing modelling approaches in novel ways.
4 Methodology
@@
In this section, we present our experimental framework for assessing Android anti-malware approaches. Section 4.1 discusses the basis for selecting studies to include. Section 4.2 presents the core set of ML models and feature selection methods that we use to extend and standardise the original studies. Section 4.3 describes the evaluation dataset, and how it was collected. Section 4.4 summarises how we extract features from Android applications through static and dynamic analysis. Section 4.5 outlines the metrics and procedures we use to assess ML models. Fig. 1 provides an overview.
4.1 Choice of studies to reimplement
There have been a large number of studies which have applied feature-based ML approaches to Android malware detection. It is not feasible, or desirable, to reimplement all of these. Instead, guided by existing reviews of the literature [Naway2018, Alqahtani2019a, Bayazit2020, Wang2020, Muzaffar2022], we have chosen to reimplement (a) seminal studies that introduced specific approaches, and (b) studies which built on these approaches in notable ways. The latter includes those which further refined the approach, considered alternative representations, or applied more complex models. These are categorised and listed in Fig. 1, in each case highlighting the features which were used to build ML models, alongside other aspects such as the use of feature selection methods and the use of deep learning models. A primary guiding principle was to select studies which explored a diversity of underlying features, in order to objectively compare their utility. A secondary principle was to include studies which used different ML models, in order to objectively gain insight into the influence of model choice and complexity.
We focus on the binary classification task of discriminating malware from benign applications. However, some studies (e.g. [son2021risk, son2022risk]) consider multi-class problems, typically with the aim of discriminating between different malware families. We chose to exclude these for several reasons. First, a key aim of this study is to understand the role of different models and features in the malware detection process, and binary classification provides the most generalisable insights into this. Second, there is no universally-accepted method for labeling malware families, with the ever-changing and diverse nature of malware leading to varying taxonomies and classifications across the cybersecurity community. Again, this poses difficulties when trying to reach generalisable insights. Third, there is a significant imbalance within the natural distribution of malware types. For instance, in the dataset collected for this study, adware comprises around 70% of the samples. This introduces a challenge of handling imbalanced data, but also the challenge of separating the effects of mechanisms used to address imbalance from the effects of models and features.
Apart from correcting any clear errors of practice, during the process of reimplementation we have preserved the modelling processes as described in the original studies. All the studies are then compared within the same experimental scenario, so that we have a level playing field from which we are able to make robust insights. This includes standardising the data used to train and evaluate models (see Section 4.3), carrying out consistent hyperparameter tuning to avoid unfairness (Section 4.2), and (where appropriate) including relevant modelling combinations that were not considered in the original study in order to provide a more complete picture of its potential.
4.2 Inclusion of core ML models and feature selection algorithms
Previous studies have used a diverse range of ML models. In general, we reimplement the model(s) used in the original study wherever possible. However, to provide better consistency and comparability across studies, we also augment the original studies with the following core set of ML models, if they are not already included: support vector machine (SVM), random forest (RF), decision tree (DT) and -nearest neighbour (kNN) [hastie2009elements]. All of these are widely used both within the Android malware detection literature (as surveyed in [Muzaffar2022]) and the broader ML literature. We use grid search to tune the hyperparameters of all the models implemented. This is to ensure fairness when making comparisons between models [Lones2021].
4.3 Dataset collection
@@
We collected111Dataset available at https://researchportal.hw.ac.uk/en/datasets/android-dataset-for-malware-detection Android applications released from 2019-2021 from the stores and repositories which are typically used by end users. The lack of availability of scripts to build datasets prompted us to build crawlers to download these applications. Through periodic use, these platform-independent Python scripts can be used to maintain an up-to-date dataset. The stores we targeted were UpToDown [uptodown], APKMirror [apkmirror] and F-Droid [fdroidsite]. The scripts crawl the websites for all applications irrespective of the categories the applications are filed under. We downloaded a total of 62,000 benign applications. To the best of our knowledge, this is the most realistic and up to date benign dataset available right now.
@@
To collect malware, we used the most recent Android malware dataset from VirusShare [virusshare] for malware released in 2020. We also identified malware while downloading applications from the application stores. We used VirusTotal [virustotal] reports to label all the applications. To prevent false positives in the benign dataset, we only included applications with zero positive tags from anti-malware in VirusTotal reports. The final malware dataset consisted of 62,000 applications, matching the size of the benign dataset.
4.4 Feature extraction process
We developed an automated analysis tool, DroidDissector222Code and scripts available at https://github.com/alisolehria/DroidDissector-A-Static-and-Dynamic-Analysis-Tool-for-Android-Malware-Detection [muzaffar2023droiddissector], to extract the static and dynamic features needed for reimplementing studies. The tool’s architecture is depicted in Figure 2. It comprises two sub-systems: one for extracting static features, and another for extracting dynamic features. Full details of how these are implemented together with the features they extract are provided in A. For static analysis, the tool extracts features that capture permissions, API calls, opcodes, hardware and app components, intents and network addresses. It also extracts the API call graph, and provides features summarising use of restricted and suspicious API calls, as used in the Drebin feature set. For dynamic analysis, the tool extracts features that capture system and API call usage, alongside network traffic and system logs. To provide scalability of large datasets, dynamic analysis is done whilst the application is running in an emulator. Unlike other commonly-used tools for dynamic analysis, DroidDissector runs on current versions of Android.
4.5 Model evaluation procedure